# Machines are more productive than humans until they aren’t, and vice versa

### v2

@@ Abstract
With the growth of artificial skills, organizations may increasingly confront with the problem of optimizing skill policy decisions guided by economic principles. This paper addresses the underlying complexity of this challenge by developing an in-silico framework based on Monte Carlo simulations grounded in empirical realism to analyze the economic impact of human and machine skills, individually or jointly deployed, in the execution of tasks presenting varying levels of complexity. Our results provide quantitative support for the established notions that automation tends to be the most economically-effective strategy for tasks characterized by low-to-medium generalization difficulty, while automation may struggle to match the economic utility of human skills in more complex scenarios. Critically, our simulations highlight that combining human and machine skills can be the most effective strategy when a high level of generalization is required, but only if genuine augmentation is achieved. In contrast, when failing to realize this synergy, the human-machine policy is severely penalized by the inherent costs of its dual skill structure, causing it to destroy value and becoming the worst choice from an economic perspective. The takeaway for decision-makers is unambiguous: in contexts requiring high generalization capabilities, simply allocating human and machine skills to a task is insufficient, and a human-machine skill policy is neither a silver-bullet solution nor a low-risk compromise. Rather, it is a critical opportunity to boost competitiveness that demands a strong organizational commitment to enabling augmentation. Also, our findings show that improving the cost-effectiveness of machine skills over time, while useful, does not replace the fundamental need to focus on achieving augmentation. By synthesizing task-based economic theory with the concept of generalization difficulty, our work offers a scalable and low-cost methodology for organizations to forecast the economic implications of their choices. This provides a foundation for testing skill policies ex ante and for scaling experiments to study more general policy implications as candidates for future empirical validation.
@@ 6.2.2 Results
Figure 9: Visualization of the distributions $\mu(\Lambda^{s})$, $\mu(\Lambda_{c}^{s})$ and $\mu(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 10: Visualization of the distributions $\sigma(\Lambda^{s})$, $\sigma(\Lambda_{c}^{s})$ and $\sigma(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
@@ 7.1 Data analysis
To understand the impact of skill policy $HM$ overall, and in conjunction with the augmenting method of preference, we break down the sets $\Lambda^{s}$, $\Lambda_{c}^{s}$ and $\Lambda_{c,d}^{s}$ by the values of the function $a$. Then, for convenience and notational clarity, we define an index called HMG, denoted as $\text{HMG}_{c_{0},d,a_{0}}^{\omega,\Lambda}$, whose role is to characterize the relative gain (or loss, if negative) achieved by the skill policy $HM$ in conjunction with the function $a=a_{0}\in\{mean,collaborate,superpower\}$ with respect to skill policy $c_{0}\in\{H,M\}$, based on the mean values of the distributions $\omega(\Lambda^{s})$, $\omega(\Lambda_{c}^{s})$ and $\omega(\Lambda_{c,d}^{s})$.
More formally:
@@ 7.2 Meta-modeling
To understand more about the importance of our variables on performance and utility, we leverage the potential of introspection analysis on meta-models (more specifically, surrogate models) learned on our synthetic data by means of i) Permutation Importance analysis and ii) SHapley Additive exPlanations method (SHAP).
@@
Since the models’ performance reveals to be capturing the signal in the data, we move forward with introspection analyses. The results of Permutation Importance, summarized in Table 13, and the SHAP swarm plots represented in Figures 15, 16, 17 and 18 show a remarkably consistent response of the two methods, allowing consideration.
@@ 7.3 Limitations
Acknowledging these limitations is pivotal to interpreting the results and to plan future developments.
@@ 8 Conclusions
With this work, we intended to offer a contribution to the economic-minded decision-makers confronting with the problem of skill policy decisions in the context of cognitive tasks characterized by a spectrum of generalization difficulty.
More specifically, we addressed two challenges: first, the estimation of the quantitative economic impact of the options of choice when navigating relatively small search spaces, leveraging a limited number of situations and simulations to produce distributions of synthetic data that can be quickly analyzed; second, in the context of larger search spaces representing higher levels of uncertainty, designing experiments comprised of a large number of simulations, whose result can allow to find more general insights and policy implications, or recommendations.
In this frame of reference, our work develops by considering task-execution performance a non-deterministic outcome of complex interactions between skills and uncertain task content, crystallizing the idea of measuring the economic impact of skill policy decisions from the bottom-up. By combining the provisions of the seminal framework illustrated in [Farboodi and Veldkamp, 2021] with the concepts presented in [Chollet, 2019], we sketch a model and a Monte Carlo simulation environment to generate synthetic representations of the economic impact of alternative skill policies involving human and/or machine skills confronting with varying levels of generalization difficulty.
The simulation’s setup conditions are set to establish a clear baseline, with Beta-distributed probabilities modeling a scenario where the state of technological advancement for machine skills is not yet able to provide consistent and widespread performance when confronting with tasks requiring high levels of generalization difficulty. More specifically, initially, the machine skill is superior on low-to-medium difficulty tasks, reflecting its high initial efficiency. Over time, the human skill evolves, partially closing the performance gap on low and medium-difficulty tasks, while simultaneously demonstrating a widening advantage on high-difficulty tasks. The margin of contribution for automation is set to be empirically higher than for human labor to reflect lower variable costs and higher economies of scale. Human-machine interaction, initially floating in a range compatible with the sum of the costs of both human and machine skills, improves its margin of contribution over epochs up to a factor of 2x. No constraint is set for the combinations of the minimum performance triggering an error shock and its unitary cost. This setup, combined with non-linear performance-to-output relationships, contribute to characterize an experimental scenario that is grounded in empirical pragmatism, suitable to representing a common real-world case.
@@
Next, by learning Random Forest and Gradient Boosting regression meta-models on our experimental synthetic data, and conducting introspection analyses through Permutation Importance tests and SHAP, we get additional support to the idea that skill policy decisions combined with the design of effective augmentation are the important predictors for economic outcomes within the proposed simulation environment and setup. Also important are the the factors conditioning the emergence of economic consequences of errors, and the baseline margin of contribution associated with skill policies based on human-machine interaction.
@@
The key takeaway for organizational leaders and policymakers is that skill policy should not be limited to the traditional binary choice of full automation or human-only labor. Our findings suggest that a human-machine integration policy presents a promising third option. However, it’s crucial to note that this approach only yields benefits when it creates genuine augmenting effects. Without such synergy, the inherent costs of the human-machine model can make it less efficient, and the increasing costs due to errors can drag the net economic effect down to zero, or even in a negative danger zone. Therefore, while human-machine integration is a significant opportunity, it is not a magic button but rather a strategic choice that requires full commitment to generate a competitive advantage.
Moreover, our results quantitatively support the notions that, while automation is most beneficial for low-to-medium difficulty tasks, it may be insufficient for high-difficulty scenarios where the cost of machine errors can erode all economic utility generated by high unitary margins of contribution. This conclusion is contingent upon the initial assumptions pertaining to the probability distributions of performance for human and machine skills. In our framework, machine skills are not omniscient, and high generalization difficulty represents their fundamental limitation. Should advancements successfully address this Achilles heel, the economic rationale against automating high-difficulty tasks may dissolve, opening new frontiers for value creation.
@@
Also, a second key takeaway for decision-makers pertains the importance of knowing as better as possible the process of interest, and the characteristics of the environment in which the task unfolds. Being able to estimate the characteristics of the probability distribution of task difficulty can be an advantage in the making of skill policy decisions that can affect the cost structure of the organization.
@@
By modeling the skill policy decision problem through an in-silico simulation environment able to replicate stylized facts already illustrated in the literature, we provide a promising and scalable framework for organizations to test alternative skill policies and their potential economic outcomes without incurring real-world costs. The ability to model and forecast the economic impact of skill policy options in advance can be, itself, a competitive advantage, enabling organizations to make more informed, efficient, and profitable strategic decisions in an era of unprecedented technological change.
@@ 10 Acknowledgments
We thank Prof. Enrico Minelli and Prof. Carmen Marchiori for the stimulating discussion about data economy in the early phases of this study. AI tools have been used for light copyediting on small portions of this manuscript. All errors are our own. No conflict of interest involving the author exist at the time of this draft is published.
@@ A.2 Results, additional material
Table 14: Descriptive statistics for distributions $\omega(\Lambda_{c,d}^{s})$ with detail by skill policy $c$, human-machine interaction mode $a$, and task difficulty $d$.
Figure 19: Visualization of the distributions $\rho(\Lambda^{s})$, $\rho(\Lambda_{c}^{s})$ and $\rho(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 20: Example of visualization: distributions $IQR(\Theta_{(c,d)}^{s})$, $IQR(Y_{(c,d)}^{s})$, $IQR(V_{(c,d)}^{s})$, $\mu(ERR_{(c,d)}^{s})$ and $IQR(U_{(c,d)}^{s})$.
Figure 21: Example of visualization: distributions $SK(\Theta_{(c,d)}^{s})$, $SK(Y_{(c,d)}^{s})$, $SK(V_{(c,d)}^{s})$, $\mu(ERR_{(c,d)}^{s})$ and $IQR(U_{(c,d)}^{s})$.
@@ A.3 EDA, additional material
Figure 22: Analysis of individual simulations (1/7)
Figure 23: Analysis of individual simulations (2/7)
Figure 24: Analysis of individual simulations (3/7)
Figure 25: Analysis of individual simulations (4/7)
Figure 26: Analysis of individual simulations (5/7)
Figure 27: Analysis of individual simulations (6/7)
Figure 28: Analysis of individual simulations (7/7)
@@ A.4.1 A practical example
In a typical fashion, the predictive performance of the models and the corresponding economic impact are estimated for human-exclusive of machine-exclusive skill policy options without distinction by task difficulty (this is the measure captured by the metrics accompanying the models, as described in the Context). In Table 16 we illustrate the situation characterizing our example,with the models’ performance the related total cost. The total cost sums the cost of errors and the cost for development and operation. The total expected cost of error is computed multiplying the MAE, estimated through the MAPE, by the individual cost of error in percentage (10%, as specified in the Context). The incremental cost of skills is, for each period, the sum of the development cost and operational cost. For simplicity, the total cost is not actualized and the overall period of interest is 3 years. The optimization objective is to minimizing the total cost.
@@
Then, in Table 17 we see an example of economic impact of the same human and machine skills selectively trained and tested with low-difficulty examples (city area) and high-difficulty examples (new area). By selectively deploying full automation through Model B on predictions expected to require low generalization difficulty, and the human-machine skill policy to high-difficulty predictions combining the skills of Junior sales reps with Model B, we achieve a total cost of 440 + 695 = 1,135, reducing of an additional 26.8% the total cost with respect to the H_Senior skill. Crucially, while the combinations A+Senior and B+Senior provide no augmentation, the combinations A+Junior and B+Junior present a visible and strong augmenting effect respectively, thus containing the cost of errors despite of the operational costs.

### v3

@@ Abstract
With the growth of artificial skills, organizations are increasingly confronting with the problem of optimizing skill policy decisions guided by economic principles. This paper addresses the underlying complexity of this challenge by developing an in-silico framework based on Monte Carlo simulations grounded in empirical realism to analyze the economic impact of human and machine skills, individually or jointly deployed, in the execution of tasks presenting varying levels of complexity. Our results provide quantitative support for the established notions that automation tends to be the most economically-effective strategy for tasks characterized by low-to-medium generalization difficulty, while automation may struggle to match the economic utility of human skills in more complex scenarios. Critically, our simulations highlight that, when high level of generalization is required and the cost of errors is high, combining human and machine skills can be the most effective strategy, but only if genuine augmentation is achieved. In contrast, when failing to realize this synergy, the human-machine policy is severely penalized by the inherent costs of its dual skill structure, causing it to destroy value and becoming the worst choice from an economic perspective. The takeaway for decision-makers is unambiguous: in complex and critical contexts, simply allocating human and machine skills to a task may be insufficient, and a human-machine skill policy is neither a silver-bullet solution nor a low-risk compromise. Rather, it is a critical opportunity to boost competitiveness that demands a strong organizational commitment to enabling augmentation. Also, our findings show that improving the cost-effectiveness of machine skills over time, while useful, does not replace the fundamental need to focus on achieving augmentation. By synthesizing task-based economic theory with the concept of generalization difficulty, our work offers a scalable and low-cost methodology for organizations to forecast the economic implications of their choices. This provides a foundation for testing skill policies ex ante and for scaling experiments to study more general policy implications as candidates for future empirical validation.
@@ 6.2.2 Results
Figure 9: Main experiment, visualization of the distributions $\mu(\Lambda^{s})$, $\mu(\Lambda_{c}^{s})$ and $\mu(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 10: Main experiment, visualization of the distributions $\sigma(\Lambda^{s})$, $\sigma(\Lambda_{c}^{s})$ and $\sigma(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
@@ 7.1 Data analysis
To understand the impact of skill policy $HM$ overall, and in conjunction with the augmenting method of preference, we break down the sets $\Lambda^{s}$, $\Lambda_{c}^{s}$ and $\Lambda_{c,d}^{s}$ by the values of the function $a$. Then, for convenience and notational clarity, we define an index called HMG, denoted as $\text{HMG}_{c_{0},d,a_{0}}^{\omega,\Lambda}$, whose role is to characterize the relative gain (or loss, if negative) achieved by the skill policy $HM$ in conjunction with the function $a=a_{0}\in\{mean,collaborate,superpower\}$ with respect to skill policy $c_{0}\in\{H,M\}$, based on the mean values of the distributions $\omega(\Lambda^{s})$, $\omega(\Lambda_{c}^{s})$ and $\omega(\Lambda_{c,d}^{s})$. More formally:
@@
Figure 14: Distributions of the values in the data sets $\mu(U^{s})$, $\mu(U^{s}_{c})$ and $\mu(U^{s}_{c,d})$, grouped by skill policy ($c$) and task difficulty ($d$)
@@
As our experimental results are grounded on a setup for the distributions $P(\theta)_{c,d,e}$ representing a relatively low performance of the machine-exclusive skill $M$ in high-difficulty tasks, we stress-test our findings by running the same experiment with an alternative setup for $P(\theta)_{c={M},d,e}$ -see Figure 15. A visualization of the distributions of the data resulting from the stress test is presented in Appendix, Section A.2, Figure 21 and Figure 22.
Figure 15: Distribution $P(\theta)_{c={M},d,e}$ used for the core experiment (left), vs the distribution $P(\theta)_{c={M},d,e}$ used in the stress-test experiment (right). As the reader would promptly notice, the improvement assumed for the machine skill by the stress-test setup is remarkable.
By means of a preliminary verification step, we first check that the values $\overline{\mu(U^{s}_{c_{0}=H,d})}$ in Table 6 and Table 12 are the same, as they must be; second, we check that the values of $\overline{\mu(U^{s}_{c_{0}=M,d})}$ and $\overline{\mu(U^{s}_{c_{0}=HM,d})},\forall a$ are increased with respect to the core experiment, as they intuitively should. This result is an obvious consequence of the new setup, and deserves no particular comment. Instead, it is more interesting to look at Tables 13 and 15, where we can appreciate the impact of the stress-test on central tendency and dispersion of utility over the simulations.
Table 12: Stress test with high machine skills, Analysis of $HMG_{c_{0},d,a_{0}}^{\mu,\Theta}$
The data presented in Table 13 suggests that the skill policy $M$ now becomes the single most attractive option from the economical perspective across all task difficulties, with no augmenting effects being able to close the gap between the human-machine skill policy versus the machine-exclusive one. However, as detailed in Table 14 and illustrated in Figure 16, focusing the lens on situations characterized by potentially high cost of error, the human-machine skill policy $HM$ featuring augmenting effects still prevails in terms of utility over the machine exclusive policy $M$ when the task difficulty is high. This reinforces the idea that, whatever the performance achieved by artificial skills, the factors determining the occurrence of errors and the magnitude of their consequences (in our case, quantified by economic metrics) impose a cap to the degree of automation that can be achieved by a responsible agent.
Table 13: Stress test with high machine skills, Analysis of $HMG_{c_{0},d,a_{0}}^{\mu,U}$
Table 14: Stress test with high machine skills, Analysis of $HMG_{c_{0},d,a_{0}}^{\mu,U}$ in conditions characterized by potential high cost of errors
Figure 16: Distributions of the values in the data sets $\mu(U^{s})$, $\mu(U^{s}_{c})$ and $\mu(U^{s}_{c,d})$, grouped by skill policy ($c$) and task difficulty ($d$), in our stress test, with focus on situations characterized by high cost of errors ($t_{err}\leq 0.1$ and $c_{err}\geq 0.80$)
Finally, a review of the Table 15 confirms that the $HM$ policy tends to reduce the dispersion of utility values with respect to skill policies $H$ and $M$, with augmentation strengthening this effect.
Table 15: Stress test with high machine skills, Analysis of $HMG_{c_{0},d,a_{0}}^{\sigma,U}$
@@ 7.2 Meta-modeling
To understand more about the importance of the model’s variables on performance and utility, we leverage the potential of introspection analysis on meta-models (more specifically, surrogate models) learned on our synthetic data by means of i) Permutation Importance analysis and ii) SHapley Additive exPlanations method (SHAP).
@@
Since the models’ performance reveals to be capturing the signal in the data, we move forward with introspection analyses. The results of Permutation Importance, summarized in Table 17, and the SHAP swarm plots represented in Figures 17, 18, 19 and 20 show a remarkably consistent response of the two methods, allowing consideration.
@@ 8 Conclusions
With this work, we intend to offer a contribution to the economic-minded decision-makers confronting with the problem of skill policy decisions in the context of cognitive tasks characterized by a spectrum of generalization difficulty.
More specifically, we address two challenges: first, the estimation of the quantitative economic impact of the options of choice when navigating relatively small search spaces, leveraging a limited number of situations and simulations to produce distributions of synthetic data that can be quickly analyzed; second, in the context of larger search spaces representing higher levels of uncertainty, the design of experiments involving a large number of simulations whose result can allow to find more general insights and policy implications.
In this frame of reference, we model task execution performance as a non-deterministic outcome of complex interactions between skills and uncertain task content, crystallizing the idea of measuring the economic impact of skill policy decisions from the bottom-up. By combining the provisions of the seminal framework illustrated in [Farboodi and Veldkamp, 2021] with the concepts presented in [Chollet, 2019], we define a theoretical model and a practical Monte Carlo simulation environment to generate synthetic representations of the economic impact of alternative skill policies involving human and/or machine skills confronting with varying levels of generalization difficulty.
Our experimental setup establishes a baseline characterizing a specific stage of technological maturity, using Beta distributions to model machine skills that excel at tasks of low-to-medium difficulty while being less effective on complex tasks demanding high levels of generalization. This performance profile for machine skills, which are set to be static once trained, is combined with human skills that evolve over epochs, partially closing the performance gap with machine skills on low and medium-difficulty tasks, and widening their advantage on high-difficulty tasks. The margin of contribution for automation is set to be empirically higher than for human labor to reflect lower variable costs and higher economies of scale. Human-machine interaction, initially floating in a range compatible with the sum of the costs of both human and machine skills, improves its margin of contribution over epochs up to a factor of 2x. No constraint is set for the combinations of the minimum performance triggering an error shock and its unitary cost. This setup, combined with non-linear performance-to-output relationships, contribute to characterize an experimental scenario that is grounded in empirical pragmatism, suitable to representing a common real-world case.
@@
A second experiment, aimed at stress-testing our findings in a setup characterized by machine skills achieving high performance across all task difficulties, confirms that automation can progressively extend its economic effectiveness to a wider set of situations as the machine skills improve. However, the human-machine skill policy still tends to prevail in high-difficulty task when the cost of error is high. Also, both our experiments show that the combination of human and machine skills tends to reduce the dispersion of values for utility, and that this reduction is more evident when augmenting effects are involved.
@@
Finally, we learn Random Forest and Gradient Boosting regression meta-models on our core experimental data to further analyze the relationship between the model variables and responses. To this end, we deploy Permutation Importance tests and SHAP. The results of the introspective analysis provide additional support to the idea that task difficulty, skill policy and the level of augmentation achieved are the most important predictors for economic impact within the proposed simulation environment and setup. Also important are the the factors conditioning the emergence of economic consequences of errors, and the baseline margin of contribution associated with skill policies based on human-machine interaction.
@@
Then, the primary takeaway for organizational leaders and policymakers is that in tasks characterize by high difficulty and high cost of error, a human-machine skill policy delivering strong augmenting effects represents an opportunity to achieve better economic results with respect to the human-exclusive or machine-exclusive options. However, without augmenting effects, the inherent costs of the human-machine model can make it less efficient, and the increasing costs due to errors can drag the net economic impact of the dual-skill policy down to zero, or even in a negative danger zone. As a consequence, a strategy based on human-machine integration is not a magic button nor a risk-mitigating solution, but a strategic entrepreneurial option of choice that requires full commitment to generate a competitive advantage.
At the same time, our core results quantitatively support the established notion that automation tends to be the most effective choice for low-to-medium difficulty tasks, with a progressive expansion of the boundaries of economical effectiveness as machine skills improve their performance in task characterized by medium-to-high difficulty. However, when tasks combine high difficulty with high cost of errors, a human-machine skill policy carefully designed for augmentation still represents the most effective option for the responsible and economic-minded agent.
@@
A second-level takeaway for decision-makers pertains the importance of knowing as better as possible the process of interest, and the characteristics of the environment in which the task unfolds. More specifically, being able to obtain high-quality estimates of the probability distribution of difficulty for the task of interest, together with robust assessments of the performance of the candidate skills by task difficulty, can allow to leverage the potential of simulation towards better supporting skill policy decisions.
@@
By modeling the skill policy decision problem through an in-silico simulation environment able to replicate stylized facts already established in the literature, we provide a promising and scalable framework for organizations to test alternative skill policies and their potential economic outcomes without incurring real-world costs. The ability to model and forecast the economic impact of skill policy options in advance can be, itself, a competitive advantage, enabling organizations to make more informed, efficient, and profitable strategic decisions in an era of unprecedented technological change.
@@ 10 Acknowledgments
We thank Prof. Enrico Minelli and Prof. Carmen Marchiori for the stimulating discussion about data economy in the early phases of this study. AI tools have been used for light copyediting on small portions of this manuscript. All errors are our own. No conflict of interest involving the author exist at the time this draft is published.
@@ A.2 Results, additional material
Table 18: Main experiment, visualization of the distributions $\omega(\Lambda_{c,d}^{s})$ with detail by skill policy $c$, human-machine interaction mode $a$, and task difficulty $d$.
Figure 21: Stress-test experiment, visualization of the distributions $\mu(\Lambda^{s})$, $\mu(\Lambda_{c}^{s})$ and $\mu(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 22: Stress-test experiment, visualization of the distributions $\sigma(\Lambda^{s})$, $\sigma(\Lambda_{c}^{s})$ and $\sigma(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 23: Main experiment, visualization of the distributions $\rho(\Lambda^{s})$, $\rho(\Lambda_{c}^{s})$ and $\rho(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 24: Main experiment, visualization of the distributions $IQR(\Theta_{(c,d)}^{s})$, $IQR(Y_{(c,d)}^{s})$, $IQR(V_{(c,d)}^{s})$, $\mu(ERR_{(c,d)}^{s})$ and $IQR(U_{(c,d)}^{s})$.
Figure 25: Main experiment, visualization of the distributions $SK(\Theta_{(c,d)}^{s})$, $SK(Y_{(c,d)}^{s})$, $SK(V_{(c,d)}^{s})$, $\mu(ERR_{(c,d)}^{s})$ and $IQR(U_{(c,d)}^{s})$.
@@ A.3 EDA, additional material
Figure 26: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 1/7)
Figure 27: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 2/7)
Figure 28: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 3/7)
Figure 29: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 4/7)
Figure 30: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 5/7)
Figure 31: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 6/7)
Figure 32: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 7/7)
@@ A.4.1 A practical example
In a typical fashion, the predictive performance of the models and the corresponding economic impact are estimated for human-exclusive and machine-exclusive skill policy options without distinction by task difficulty (this is the measure captured by the metrics accompanying the models, as described in the Context). In Table 20 we illustrate the situation characterizing our example,with the models’ performance the related total cost. The total cost sums the cost of errors and the cost for development and operation. The total expected cost of error is computed multiplying the MAE, estimated through the MAPE, by the individual cost of error in percentage (10%, as specified in the Context). The incremental cost of skills is, for each period, the sum of the development cost and operational cost. For simplicity, the total cost is not actualized and the overall period of interest is 3 years. The optimization objective is to minimizing the total cost.
@@
Then, in Table 21 we see an example of the economic impact of the human and machines skills as seen through the specific lens of low-difficulty examples (properties in the typical area served by the agency) and high-difficulty examples (properties in the new areas). For machine skills, Model A and Model B are selectively trained and tested on low-difficulty and high-difficulty examples, creating at all effects four distinct models (as illustrated in the previous section, the test phase may require to consider multiple perturbations of the training/test set composition, but the final model is trained on all the examples available). For human skills it is not practically possible to compartmentalize the training phase on specific examples, so the models H_Senior and H_Junior are simply selectively tested on low/high-difficulty sets to determine more specific performance assessments with respect to Table 20.
Then, selectively deploying Model B (automation) on predictions expected to require low generalization difficulty, and the human-machine skill policy to high-difficulty predictions combining the skills of Junior sales reps with Model B, a total cost of 440 + 695 = 1,135 is achieved, reducing it of an additional 26.8% with respect to the H_Senior skill deployed as one-fits-all solution.
Crucially, while the combinations A+Senior provide low augmentation and B+Senior a moderate one, the combinations A+Junior and B+Junior present a visible and strong augmenting effect respectively, thus containing the cost of errors despite of the operational costs. As the data is not empirically gathered and constructed only for illustrative purposes, we cannot analyze from the bottom-up the specific reasons why these augmenting effects emerge. However, intuitively, this may happen when the models are extremely diverse in their ability to capture the nature of signals present in the data, thus making one model very effective in achieving high performance when the other systematically fails, and vice versa. In this case, the logic of interaction and cooperation between the models towards determining the final response is a critical factor for success, deserving a specific effort (for example, in a first instance, the relatively simple weighting of the responses provided by the models part of the $HM$ ensemble can be learned on data, after a reasonable period of dedicated data collection; in a further and more refined step of development, one may also think to find additional variables that allow to set specific weights for specific situations, thus creating a set of expert ensembles based upon standard elementary models).